OpenStudy (anonymous):

no +ve integer x & y such that, x^2 - 3xy + 2y^2 = 10

OpenStudy (agent0smith):

not sure what the exact question is, but it seems like you need to factor it:\[\large (x -2y)(x-y) = 10\]now each of the brackets must be factors of 10 (since i'm assuming they're only allowing integers)

OpenStudy (agent0smith):

eg one bracket must be equal to 1, the other must be equal to 10 or one is equal to 2, and the other 5 those are the only factors of 10. You can use them to make simultaneous equations.

OpenStudy (anonymous):

Prove or disprove there are no positive integers x and y such that; \[x^2 -3xy+2y^2=10\]

OpenStudy (anonymous):

Not sure how to do it through induction.

OpenStudy (agent0smith):

x-2y = 10 and x-y = 1 or x-2y = 1 and x-y = 10 or x-2y = 5 and x-y = 2 or x-2y = 2 and x-y = 5 with those you can prove it and it won't take long.

OpenStudy (agent0smith):

If there are any positive integers, those will find them.

OpenStudy (anonymous):

cheers mate

OpenStudy (agent0smith):

oh, actually, you'd also need to check for cases when both factors are -ve... ie -1 and -10, and -2 and -5